A hydraulic press consist of two cylinders of cross-sectional area 0.2m square and 5.0m square. The piston in the smaller cylinder is pushed down with a force of 100N through a distance of 0.2m. Calculate the pressure transmitted in fluid.
1
Expert's answer
2018-09-07T16:18:08-0400
The pressure transmitted in fluid:
P=p_2-p_1=F/A_2 -F/A_1 =100/0.2-100/5=480 Pa.